# Chain of Summaries: Summarization Through Iterative Questioning

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly using external web content. However, much of this content is not easily digestible by LLMs due to LLM-unfriendly formats and limitations of context length. To address this issue, we propose a method for generating general-purpose, information-dense summaries that act as plain-text repositories of web content. Inspired by Hegel’s dialectical method, our approach, denoted as Chain of Summaries (CoS), iteratively refines an initial summary (thesis) by identifying its limitations through questioning (antithesis), leading to a general-purpose summary (synthesis) that can satisfy current and anticipate future information needs. Experiments on the TriviaQA, TruthfulQA, and SQUAD datasets demonstrate that CoS outperforms zero-shot LLM baselines by up to 66% and specialized summarization methods such as BRIO and PEGASUS by up to 27%. CoS-generated summaries yield higher Q&A performance compared to the source content, while requiring substantially fewer tokens and being agnostic to the specific downstream LLM. CoS thus resembles an appealing option for website maintainers to make their content more accessible for LLMs, while retaining possibilities for human oversight.
Chain of Summaries: Summarization Through Iterative Questioning
William Brach Slovak Technical University william.brach@stuba.sk          Lukas Galke Poech University of Southern Denmark galke@imada.sdu.dk
1 Introduction
@@
Figure 1: Chain of Summaries (CoS) method implements Hegel’s dialectical method for document summarization, iteratively refining an initial summary (thesis) through challenging questions (antithesis) to produce an information-dense optimal summary (synthesis) capable of answering diverse queries.
@@
Large language models (LLMs) are given increasing autonomy to navigate and process information on the Web (e.g., Search and Deep Research functionalities in commercial systems). However, most of the web content is not easy to digest for LLMs and challenges their context size, leading to necessary compromises in performance and potential instabilities. It has been proposed to make the content on the web more accessible to LLMs by running parallel versions of the website in an LLM-friendly markdown format howard2024llmstxt. But how can we turn the plethora of existing content on the Web into information-dense chunks of plain text that capture the essence of the original content along various, potentially unpredictable, aspects?
@@
Here, we look at this problem from the angle of summarization – in the sense of making information better available for LLMs to consume. Concretely, if we had a robust summarization method that captured the essence of the content in various aspects, we could hope that a variety of future questions could be answered just based on the summary. If questions come up that cannot be answered from the summary, the full website could be consulted, and the summary could be revised accordingly for future requests. This way, the plain-text summary would effectively act as a cache for websites, which can be implemented both server-side (e.g., website maintainers) and client-side (e.g., Deep Research systems) and can be subject to human oversight as needed. The key idea of this paper is to simulate this process through an instantiation of Hegel’s dialectical method sep-hegel-dialectics with LLMs. We hypothesize that an initial summary draft (the thesis in Hegel’s terms) can be iteratively refined based on its capability to answer synthetically generated questions (the antithesis), ultimately crystallizing a general-purpose summary (synthesis) that facilitates answering as many unseen future questions as possible (cache hit, in the analogy). Representation of the summary must be in natural language for two reasons. First, it ensures human verifiability; website maintainers must be able to easily read and validate the information that LLM agents will consume, preventing the spread of inaccurate or outdated content. Second, a plain-text format is model-agnostic, creating a universal and durable representation that any LLM agent can use, in contrast to model-specific formats like dense embeddings. We denote this procedure as the Chain of Summaries (CoS; Figure 1).
@@
To test the effectiveness of CoS, we focus on the experiments with question-answering datasets because they allow us to take questions from the dataset that we can run against the summary of the content. In addition, QA datasets enable us to compare how helpful synthetic questions are compared to questions taken from the dataset of these datasets (human-crafted). To achieve representative selection, we run experiments with a large commercial model (GPT-4o-mini openai2024gpt4omini) as well as publicly available models of smaller sizes: Llama-3.2:3B grattafiori2024llama3herdmodels and Qwen-2.5:7B qwen2.5 to perform the roles in the CoS approach: generate the initial summary, synthesize questions, determine whether they can be answered from the summary, and iteratively refine the summary. Although technically, models can be mixed and matched, we keep the LLM backbone constant for our experiments. It is expected that GPT-4o-mini will perform best, yet we are interested in how much performance can be retained using smaller, local models. We further experimented with several different hyperparameters of the CoS approach and found that iteratively refining the summary (e.g., iterate 10 times with 1 question per iteration) has a clear advantage over presenting all questions at once (e.g., 50 questions). Our experiments show that this method of creating information-dense summaries not only shortens the content but also makes the core information more accessible to LLMs. It enables higher question-answering performance than using the full source document.
Finally, we investigate how much the multi-step CoS summarization method can be reduced to a single step. For this, we first fine-tune an encoder-only model to estimate the quality of downstream QA performance (through regression on the main evaluation metric). Then we use this encoder-only model as a reward model for reinforcement fine-tuning Qwen2:0.5B-Instruct qwen2 and Qwen3:0.6B qwen3 via Group Relative Policy Optimization (GRPO) shao2024deepseekmathpushinglimitsmathematical. Our results show that CoS consistently outperforms the strong zero-shot summarization baseline, given the same backbone LLM – suggesting that CoS improves summaries’ QA performance. Our reinforcement-finetuned Qwen2:0.5B-Instruct and Qwen3:0.6B models performed slightly worse than the pure CoS approach, indicating that the incremental procedure is beneficial to obtain a high-quality summary with the limitation of potential reward hacking. Similarly, CoS outperforms the established specialized summarization baselines, such as BRIO liu2022briobringingorderabstractive and PEGASUS zhang2019PEGASUS, as well as strong zero-shot summaries generated by LLM.
In sum, this paper makes three contributions:
@@
  • We fine-tune a BERT model to predict a summary performance in a downstream QA task, which we then use as a reward model for reinforcement fine-tuning (GRPO) of an LLM, testing whether the multi-step procedure can be reduced to a single step.
2 Related Work
Text summarization has seen advances from both abstractive and extractive single-pass methods. Abstractive models include PEGASUS 10.1016/j.neucom.2024.128280; zhang2023benchmarkinglargelanguagemodels, with its sentence-reconstruction pre-training, and BRIO Gambhir2017, which improves candidate summary ranking. In extractive summarization, recent work like DiffuSum zhang-etal-2023-diffusum uses diffusion models to match summary representations with source sentences. summarization via Large Language Models is becoming more and more popular, but often suffers from the "middle curse" ravaut-etal-2024-context; liu2023lostmiddlelanguagemodels, a strong positional bias where models struggle to utilize information from the middle of long contexts. Recent work has focused on improving LLM-retrieval integration through two main approaches. The first treats LLMs as black boxes while refining retrieved context: Replug shi-etal-2024-replug uses perplexity scores to train the retriever, Fit-Rag 10.1145/3676957 employs bi-label scoring for relevance and LLM preferences, and Recomp xu2023recompimprovingretrievalaugmentedlms compresses documents to reduce noise. The second approach fine-tunes LLMs for better knowledge integration. Atlas 10.5555/3648699.3648950 jointly pre-trains retriever and generator, while Raven huang2024ravenincontextlearningretrievalaugmented introduces Fusion-in-Context Learning. Ra-Dit lin2024raditretrievalaugmenteddualinstruction proposes dual instruction tuning, ChatQA liu2024chatqasurpassinggpt4conversational uses two-stage instruction tuning, and Rankrag 10.5555/3737916.3741766 unifies ranking and generation in a single model. Genki shen2025genkienhancingopendomainquestion integrates knowledge directly into model parameters via fine-tuning. To address limitations of single-pass generation, iterative refinement approaches have emerged. SELF-REFINE 10.5555/3666122.3668141 enables LLMs to generate, critique, and revise their outputs, while SummIt zhang2023summit combines summarization with self-evaluation to improve faithfulness without supervised training. SuRe kim2024sure generates multiple answer summaries and uses the summarization process itself as a verification mechanism. Building on these advances, our work introduces a question-guided iterative framework that explicitly optimizes for information density by ensuring summaries can answer the most important questions about the source text, bridging QA-based evaluation with retrieval-augmented generation and iterative refinement.
3 Chain of Summaries
@@
The Chain of Summaries (CoS) approach is inspired by Hegel’s dialectical method, whereby knowledge evolves through the stages of thesis, antithesis, and synthesis. As shown in Figure 1, our system works in phases with several interconnected components that progressively enhance the summaries’ quality. Our proposed CoS approach works as follows (see also Figure 1):
@@
Initial summary generation (Thesis): Given a source document, we first use an LLM to generate a zero-shot summary. This initial thesis provides a preliminary understanding that inevitably contains internal contradictions and limitations.
@@
Question Generation (Antithesis): Based on the original document, we generate synthetic question-answer pairs using the prompt in Appendix A. These questions act as the antithesis by challenging the summary’s claim to comprehensiveness, highlighting areas where key information is omitted.
@@
Iterative Refinement Process (Movement towards Synthesis): At the core of our approach lies a dialectical cycle that systematically resolves contradictions:
  • Evaluation Phase: The Recognition of Contradiction component tests the current summary’s ability to answer questions from training and validation sets using the prompt in Appendix A. Questions that cannot be answered correctly reveal the thesis’s limitations.
@@
  • Refinement Phase: A ’synthesis engine’ component creates a new synthesis that incorporates missing information while maintaining conciseness. Using the source document, current summary, and unanswered questions, it creates an improved representation that transcends the previous version. The prompt for this phase is displayed in Appendix A.
@@
Continued Dialectical Progression (Iteration): This evaluate-refine cycle repeats for $i$ iterations, with each cycle representing a dialectical movement towards the Absolute — a summary that maximizes information density for prospective downstream question-answering tasks. Once all iterations have been completed, the system selects the best synthesis based on its validation performance.
Hyperparameters for Chain of Summaries
@@
The movement towards synthesis can be adjusted using several hyperparameters that govern the iterative approach. Number of iterations ($i$), this number determines how many iterations the summary undergoes. Higher values allow for a more comprehensive evolution, but there is a risk of diminishing returns or over-refinement. The number of questions for refinement ($iq$) controls how many unanswered questions from each evaluation phase are used to guide the subsequent synthesis. This parameter balances focused improvement (low values) against broader coverage (high values).
4 Experimental Setup
4.1 Datasets
@@
To directly assess our primary objective of creating summaries maximally useful for downstream QA tasks, we selected standard QA datasets. By testing question-answering using only the summary, we measure information density and utility. Our evaluation uses TriviaQA 2017arXivtriviaqa, TruthfulQA lin2022truthfulqameasuringmodelsmimic, and SQUAD v1.1 rajpurkar-etal-2016-squad. To ensure a valid evaluation of our iterative approach, we select documents containing at least three test questions, providing sufficient representation to assess performance across iterations.
4.2 LLM Backbones
@@
4.3 Baselines
To provide context for the performance of our CoS approach, we compare it with a range of benchmarks designed to evaluate its effectiveness against direct LLM applications, alternative dense representations, and other summarization techniques. We use identical LLM backbones to generate a direct, one-pass summary. This is the simplest application of LLMs — zero-shot summarization — and shows how our iterative CoS process improves performance. The Full Source Content baseline represents a source document as context for the downstream QA task. This serves as a benchmark to determine whether our summarization process merely condenses information or actively makes it more accessible and useful for an LLM by mitigating issues such as the ’lost-in-the-middle’ effect, for example. BRIO and PEGASUS represent supervised abstractive summarization approaches, enabling comparison between our iterative method and specialised summarization techniques.
4.4 Evaluation Metrics
@@
Dataset-specific Evaluation Measures
5 Results
@@
Chain of summaries (GPT-4o-mini) 0.48 0.80 30.34 12.69 0.46 0.27
Zero shot summary (GPT-4o-mini) 0.43 0.73 24.04 10.67 0.41 0.27
@@
Content - 0.76 - 10 - 0.28
CoS-Finetuned-Qwen3:0.6B - 0.73 - 9.55 - 0.25
CoS-Finetuned-Qwen2:0.5B-Instruct - 0.74 - 8.77 - 0.24
@@
Chain of summaries (Llama3.2:3B) 0.39 0.62 29.31 14.77 0.34 0.20
Zero shot summary (Llama3.2:3B) 0.21 0.52 18.34 8.89 0.23 0.18
@@
Content - 0.51 - 4.52 - 0.19
CoS-Finetuned-Qwen3:0.6B - 0.50 - 5.33 - 0.19
CoS-Finetuned-Qwen2:0.5B-Instruct - 0.49 - 6.48 - 0.18
@@
Chain of summaries (Qwen2.5:7B) 0.52 0.60 46.06 10.71 0.59 0.16
Zero shot summary (Qwen2.5:7B) 0.23 0.48 18.55 7.35 0.36 0.14
@@
Content - 0.41 - 6.25 - 0.16
CoS-Finetuned-Qwen3:0.6B - 0.45 - 4.89 - 0.15
CoS-Finetuned-Qwen2:0.5B-Instruct - 0.42 - 5.33 - 0.13
Table 1: A performance comparison of summarization methods across three language model backbones on the TriviaQA, TruthfulQA, and SQUAD 1.1 datasets. Our CoS approach outperforms all baselines on generalization metrics across all model backbones and datasets. Bold values indicate the best performance, and italicized values indicate the second-best performance. The table also shows that CoS effectively retains critical information through iterative refinement.
@@
Table 1 presents the main results. CoS consistently outperforms zero-shot summarization across all models and datasets, with improvements using GPT-4o-mini as backbone on TriviaQA (0.80 vs. 0.73, a 9.6% gain) and in Llama3.2:3B on TruthfulQA (14.77 vs. 8.89, a 66% improvement). Our approach also exceeds traditional summarization methods, surpassing BRIO by 27% (0.80 vs. 0.63) and PEGASUS by 11% (0.80 vs. 0.72) on TriviaQA with GPT-4o-mini. A key finding is that CoS summaries consistently present information more effectively for downstream QA purposes than the original documents themselves. For instance, on TriviaQA with a GPT-4o-mini backbone, the CoS summary achieves a score of 0.80, which is higher than the score of 0.76 achieved using the full source content. This pattern, whereby the summaries created by CoS are more useful than the raw text, is consistent across multiple models and datasets, as demonstrated in Table 1. For memorization, the difference is particularly striking with Qwen2.5:7B on TruthfulQA, where CoS achieves 46.06 compared to just 18.55 for zero-shot summaries.
@@
Figure 2: Summary length (tokens, log scale) vs. Correct-F1 score on the TriviaQA dataset, comparing performance across model backbones and summarization methods. This visualization demonstrates CoS’s ability to create information-dense summaries that optimize the efficiency-performance trade-off across model architectures.
Summary Length
Our analysis of summary length versus generalization performance reveals distinct patterns across models within our CoS approach. As shown in Figure 2 and in more detail in Appendix B, CoS consistently outperforms traditional summarization methods while requiring substantially fewer tokens.
Using only 170 tokens, GPT-4o-mini enables CoS to achieve an F1 score of 0.80, which substantially outperforms zero-shot summarization (0.73), BRIO (0.63), and PEGASUS (0.72). The most striking comparison is with content-based approaches, which, despite achieving a competitive F1 score of 0.76, require over 11,000 tokens (full content), making them impractical for most applications. For Llama3.2:3B, CoS achieves an F1 score of 0.62, significantly exceeding both zero-shot approaches (0.52) and established methods with similar token efficiency. Qwen2.5 demonstrates the most impressive efficiency gains, with CoS achieving an F1 score of 0.60 using just 42–47 tokens, showing that our method can produce highly informative summaries in a very compact form.
@@
One of the questions for implementing our CoS approach is whether synthetic questions can effectively replace costly human-crafted ones. As shown in Table 2, synthetic questions are indeed viable. Across all three models, synthetic questions yielded comparable or superior results to human-crafted ones. With GPT-4o-mini, synthetic questions improved generalization from 0.73 to 0.80 (+0.07), slightly outperforming human-crafted questions, which improved from 0.74 to 0.79 (+0.05). Similarly, Qwen2.5:7B showed nearly identical improvements with synthetic questions (memorization: +0.29, generalization: +0.12) compared to human-crafted questions (memorization: +0.33, generalization: +0.10). The prompts for question generation and answering can be found in Appendix A.

### v2

# Chain of Summaries: Condensing Information by Anticipating  Questions

@@ Abstract.
Large Language Models (LLMs) are increasingly using external web content. However, much of this content is not easily digestible by LLMs due to LLM-unfriendly formats and limitations of context length. To address this issue, we propose a method for generating general-purpose, information-dense summaries that act as plain-text repositories of web content. Inspired by Hegel’s dialectical method, our approach, denoted as Chain of Summaries (CoS), iteratively refines an initial summary (thesis) by identifying its limitations through questioning (antithesis), leading to a general-purpose summary (synthesis) that can satisfy current and anticipate future information needs. Experiments on the TriviaQA, TruthfulQA, and SQUAD datasets demonstrate that CoS outperforms zero-shot LLM baselines by up to 66% and specialized summarization methods such as Chain of Density, BRIO and PEGASUS by up to 27%. CoS-generated summaries yield higher Q&A performance compared to the source content, while requiring substantially fewer tokens and being agnostic to the specific downstream LLM. CoS thus resembles an appealing option for website maintainers to make their content more accessible for LLMs, while retaining possibilities for human oversight.
summarization, information density, large language models, web agents
$c$s: Computing methodologies Natural language processingccs: Information systems Summarization
1. Introduction
@@
Figure 1. Chain of Summaries (CoS) method implements Hegel’s dialectical method for document summarization, iteratively refining an initial summary (thesis) through challenging questions (antithesis) to produce an information-dense optimal summary (synthesis) capable of answering diverse queries.
@@
Large language models (LLMs) are given increasing autonomy to navigate and process information on the Web (e.g., Search and Deep Research functionalities in commercial systems). However, most of the web content is not easy to digest for LLMs and challenges their context size, leading to necessary compromises in performance and potential instabilities. It has been proposed to make the content on the web more accessible to LLMs by running parallel versions of the website in an LLM-friendly markdown format (Howard, 2024). But how can we turn the plethora of existing content on the Web into information-dense chunks of plain text that capture the essence of the original content along various, potentially unpredictable, aspects?
@@
Here, we look at this problem from the angle of summarization in the sense of making information better available for LLMs to consume. Concretely, if we had a robust summarization method that captured the essence of the content in various aspects, we could hope that a variety of future questions could be answered just based on the summary. If questions come up that cannot be answered from the summary, the full website could be consulted, and the summary could be revised accordingly for future requests. This way, the plain-text summary would effectively act as a cache for websites, which can be implemented both server-side (e.g., website maintainers) and client-side (e.g., Agentic systems) and can be subject to human oversight as needed.
The key idea of this paper is to frame this process through the lens of Hegel’s dialectical method (Maybee, 2020) and implement the procedure with multiple LLM calls. Our main hypothesis is that an initial summary draft (the thesis in Hegel’s terms) can be iteratively refined based on its capability to answer synthetically generated questions (the antithesis), creating a general-purpose summary (synthesis) that facilitates answering as many unseen future questions as possible – thereby, maximizing the number of cache hits. We denote this procedure as the Chain of Summaries (CoS; Figure 1).
Representing the summary in natural language has two key advantages: First, it ensures human verifiability; website maintainers must be able to easily read and validate the information that LLM agents will consume, preventing the spread of inaccurate or outdated content. Second, a plain-text format is model-agnostic, creating a universal and durable representation that any LLM agent can use, in contrast to model-specific formats like dense embeddings from language models that are often used for retrieval-augmented generation (Lewis et al., 2020).
To test the effectiveness of our Chain of Summaries approach, we focus our experiments on question-answering datasets. This is because considering a question-answering task enables us to take questions from the dataset that we can run against the summary of the content for evaluation. This further allows us to compare how helpful synthetically generated questions are compared to questions taken from the training split of these datasets (human-crafted).
@@
With the aim of a representative selection, our experiments cover a commercial model (GPT-4o-mini (OpenAI, 2024)) as well as publicly available models of smaller sizes: Llama-3.2:3B (Grattafiori et al., 2024) and Qwen-2.5:7B (Qwen et al., 2025) to enact the different roles in the CoS framework: (i) generate the initial summary; (ii) synthesize questions; (iii) determine whether they can be answered from the summary; and (iv) iteratively refine the summary. See Figure 1 for an overview of the procedure. Although technically, models can be mixed and matched, we keep the LLM backbone constant for our experiments. It is expected that GPT-4o-mini will perform best, yet we are interested in how much performance can be retained using smaller, local models.
We further experimented with several different hyperparameters of the CoS approach and found that iteratively refining the summary (e.g., iterate 10 times with 1 question per iteration) has a clear advantage over presenting all questions at once (e.g., 50 questions). Notably, we split available questions into training, validation, and test sets, and selected the best-performing summary for each document based on validation performance before evaluating it on the test set.
A key factor of interest is to what extent the training set of questions, that is available for question-answering datasets but not for summarizing arbitrary content on the web, can be replaced by having another language model generating synthetic questions. We will show that synthetic questions are on par or almost as good in all cases.
In general, our experiments show that our method of creating information-dense summaries through generating synthetic questions and iteratively refining the summary is highly effective: It shortens the content but also makes the core information better accessible to LLMs. We find that it often leads to an even higher question-answering performance than when using the full source document as input. Chain of Summaries also outperforms competing summarization techniques, traditional ones such as BRIO and PEGASUS, or modern ones based on zero-shot prompting a language model for a summary, or other iterative approaches such as Chain of Density.
Finally, we conduct a break-even analysis to study after how many queries to an assumed website the upfront cost of the summarization procedure is amortized. We find that it takes only 13-15 queries for which the summary can be consulted instead of the full document.
In sum, this paper makes three core contributions:
@@
  • We provide a efficiency analysis showing that CoS summaries reduce per-query token consumption by over 98%, with break-even occurring after approximately 14 queries, making CoS cost-effective for repeatedly-accessed documents.
2. Related Work
Text summarization has seen advances from both abstractive and extractive single-pass methods. Abstractive models include PEGASUS (Luo et al., 2024; Zhang et al., 2023c), with its sentence-reconstruction pre-training, and BRIO (Gambhir and Gupta, 2017), which improves candidate summary ranking. In extractive summarization, recent work like DiffuSum (Zhang et al., 2023a) uses diffusion models to match summary representations with source sentences. Summarization via large language models is becoming increasingly popular, but often suffers from the ”middle curse” (Ravaut et al., 2024; Liu et al., 2023), a strong positional bias where models struggle to utilize information from the middle of long contexts.
Recent work has focused on improving LLM-retrieval integration through two main approaches. The first treats LLMs as black boxes while refining retrieved context: Replug (Shi et al., 2024) uses perplexity scores to train the retriever, Fit-Rag (Mao et al., 2025) employs bi-label scoring for relevance and LLM preferences, and Recomp (Xu et al., 2023) compresses documents to reduce noise.
The second approach fine-tunes LLMs for better knowledge integration. Atlas (Izacard et al., 2023) jointly pre-trains retriever and generator, while Raven (Huang et al., 2024) introduces Fusion-in-Context Learning. Ra-Dit (Lin et al., 2024) proposes dual instruction tuning, ChatQA (Liu et al., 2024) uses two-stage instruction tuning, and Rankrag (Yu et al., 2025) unifies ranking and generation in a single model. Genki (Shen et al., 2025) integrates knowledge directly into model parameters via fine-tuning.
To address limitations of single-pass generation, iterative refinement approaches have emerged. SELF-REFINE (Madaan et al., 2023) enables LLMs to generate, critique, and revise their outputs, while SummIt (Zhang et al., 2023b) combines summarization with self-evaluation to improve faithfulness without supervised training. SuRe (Kim et al., 2024) generates multiple answer summaries and uses the summarization process itself as a verification mechanism.
Chain of Density (Adams et al., 2023) iteratively densifies summaries while maintaining fixed length, optimizing for entity density with human preference evaluation; in contrast, our CoS approach optimizes for downstream QA performance through dialectical refinement. Building on these advances, our work introduces a question-guided iterative framework that explicitly optimizes for information density by ensuring summaries can answer the most important questions about the source text, bridging QA-based evaluation with retrieval-augmented generation and iterative refinement.
3. Chain of Summaries
@@
The Chain of Summaries (CoS) approach is inspired by Hegel’s dialectical method, whereby knowledge evolves through the stages of thesis, antithesis, and synthesis. As shown in Figure 1, our system works in phases with several interconnected components that progressively enhance the summaries’ quality.
Given a source document $D$, our method generates a set of question-answer pairs $Q=\{(q_{1},a_{1}),(q_{2},a_{2}),\ldots,(q_{n},a_{n})\}$ and partitions them into training, validation, and test sets: $Q=Q_{\text{train}}\cup Q_{\text{val}}\cup Q_{\text{test}}$. The iterative refinement process operates as follows:
@@
Initial Summary Generation (Thesis): We first use an LLM to generate a zero-shot summary $s^{0}$ from the source document $D$. This initial thesis provides a preliminary understanding that inevitably contains internal contradictions and limitations.
@@
Question Generation (Antithesis): Using the prompt in Listing 1, we generate the synthetic question-answer set $Q$ from $D$. These questions act as the antithesis by challenging the summary’s claim to comprehensiveness, highlighting areas where key information is omitted.
@@
Iterative Refinement Process (Movement towards Synthesis): For each iteration $k\in\{1,\ldots,i\}$, the dialectical cycle proceeds as:
  • Evaluation Phase: We evaluate the current summary $s^{k-1}$ against questions in $Q_{\text{train}}$ using the prompt in Listing 2. Let $U^{k}\subseteq Q_{\text{train}}$ denote the set of unanswered questions—those for which $s^{k-1}$ fails to provide correct answers. These reveal the thesis’s limitations.
@@
  • Refinement Phase: We sample a subset of $iq$ questions from $U^{k}$ and invoke a synthesis engine that takes as input the tuple $(D,s^{k-1},U^{k}_{iq})$ and produces an improved summary $s^{k}$ that incorporates the missing information while maintaining conciseness. The prompt for this phase is displayed in Listing 3.
@@
Continued Dialectical Progression (Iteration): This evaluate-refine cycle repeats for $i$ iterations, producing a sequence of candidate summaries $\mathcal{S}=\{s^{0},s^{1},\ldots,s^{i}\}$. Each iteration represents a dialectical movement towards the best summary that maximizes information density for prospective downstream question-answering tasks. The final summary $s^{*}$ is selected based on validation performance.
Summary selection based on validation performance
Since the iterative refinement process produces multiple candidate summaries (one per iteration), we employ a validation-based selection strategy to choose the final summary for each document. Specifically, we partition the available questions into training, validation, and test sets. The training questions guide the refinement process, while validation questions, never seen during refinement, are used to select the best performing summary across all iterations. Formally, for each document with candidate summaries $\mathcal{S}=\{s^{0},s^{1},\ldots,s^{i}\}$ from $i$ iterations, we select:
(1) $s^{*}=\underset{s^{k},k\in\{1,\ldots,i\}}{\operatorname{arg\,max}}\text{MeanF1}(s^{k},Q_{\text{val}})$
where $Q_{\text{val}}$ denotes the validation question set. The selected summary $s^{*}$ is then evaluated on the held-out test set $Q_{\text{test}}$ to compute the final generalization score. This procedure ensures that our reported results reflect true out of sample performance, preventing overfitting to the questions used during iterative refinement.
Hyperparameters
@@
The synthesis process is controlled by two key hyperparameters: the number of iterations $i$, which determines how many refinement cycles occur, and the number of refinement questions $iq$, which controls how many unanswered questions from $U^{k}$ guide each iteration. Higher $i$ enables more comprehensive evolution but risks over-refinement, while $iq$ balances focused improvement (low values) against broader coverage (high values).
4. Experimental Setup
4.1. Datasets
@@
To directly assess our primary objective of creating summaries maximally useful for downstream QA tasks, we selected standard QA datasets. By testing question-answering using only the summary, we measure information density and utility. Our evaluation uses TriviaQA (Joshi et al., 2017), TruthfulQA (Lin et al., 2022), and SQuAD v1.1 (Rajpurkar et al., 2016). We repurpose these datasets by extracting (document, question, answer) triples. This allows us to test if the summary retains enough information to answer the questions, replacing standard reading comprehension tests. To ensure a valid evaluation of our iterative approach, we select documents containing at least three test questions, providing sufficient representation to assess performance across iterations. All compared methods receive identical (document, question, answer) triples, ensuring fair evaluation.
4.2. LLM Backbones
@@
4.3. Baselines
To provide context for the performance of our CoS approach, we compare it with a range of benchmarks designed to evaluate its effectiveness against direct LLM applications, alternative dense representations, and other summarization techniques. We use identical LLM backbones to generate a direct, one-pass summary. This is the simplest application of LLMs — zero-shot summarization — and shows how our iterative CoS process improves performance. The Full Source Content baseline represents a source document as context for the downstream QA task. This serves as a benchmark to determine whether our summarization process merely condenses information or actively makes it more accessible and useful for an LLM by mitigating issues such as the ’lost-in-the-middle’ effect, for example. BRIO and PEGASUS represent supervised abstractive summarization approaches, enabling comparison between our iterative method and specialised summarization techniques. Chain-of-Density (Adams et al., 2023) is an iterative prompting technique that progressively increases summary information density through multiple LLM passes.
4.4. Evaluation Metrics
@@
5. Results
@@
Original Content - 0.76 - 10 - 0.28
Chain of summaries w/ GPT-4o-mini (ours) 0.48 0.80 30.34 12.69 0.46 0.27
Zero shot summary w/ GPT-4o-mini 0.43 0.73 24.04 10.67 0.41 0.27
Chain of Density - 0.73 - 9.07 - 0.26
@@
Original Content - 0.51 - 4.52 - 0.19
Chain of summaries w/ Llama3.2:3B (ours) 0.39 0.62 29.31 14.77 0.34 0.20
Zero shot summary w/ Llama3.2:3B 0.21 0.52 18.34 8.89 0.23 0.18
Chain of Density - 0.46 - 4.63 - 0.17
@@
Original Content - 0.41 - 6.25 - 0.16
Chain of summaries w/ Qwen2.5:7B (ours) 0.52 0.60 46.06 10.71 0.59 0.16
Zero shot summary w/ Qwen2.5:7B 0.23 0.48 18.55 7.35 0.36 0.14
Chain of Density - 0.36 - 5.13 - 0.12
@@
Table 1. A performance comparison of summarization methods across three language model backbones on the TriviaQA, TruthfulQA, and SQUAD 1.1 datasets. Our CoS approach outperforms all baselines on generalization metrics across all model backbones and datasets. Bold values indicate the best performance, and italicized values indicate the second-best performance. The table also shows that CoS effectively retains critical information through iterative refinement.
Zero-Shot Summary Chain of Summaries (CoS) Summary
Sherlock Holmes, created by Sir Arthur Conan Doyle, is a renowned fictional detective known for his exceptional observational and deductive skills, first appearing in 1887. Sherlock Holmes, created by Sir Arthur Conan Doyle, is a renowned fictional detective known for exceptional observation and deductive skills, first appearing in 1887.
His character has significantly influenced mystery literature and popular culture, leading to numerous adaptations and a lasting legacy as one of the most portrayed characters in film history. Holmes requires keen observation skills for deduction. Conan Doyle cited Joseph Bell as an inspiration.Holmes feels disinterested in personal relationships. A prevalent theme is Holmes’s distrust of women. He sometimes uses charm to influence others. The original setting for most investigations is Victorian London.
Robert Downey Jr. starred in adaptations directed by Guy Ritchie. Holmes’s estimated birth year is 1854 based on ”His Last Bow.” The first series is The Adventures of Sherlock Holmes. He fails to recognize unimportant facts. Holmes achieved a university degree before detection. ”The Hound of the Baskervilles” features Holmes’s return after death. He employs abductive reasoning to solve cases.
Table 2. Comparison of Zero-Shot vs. CoS Summarization Density. The Zero-Shot summary captures the global context but misses 85% of the specific factual entities. The CoS summary maintains the global context while surfacing granular details (biographical dates, inspirations, and specific adaptations) that are critical for high-fidelity information retrieval.
@@
Table 1 presents the main results, and Table 2 highlights a representative CoS summary versus a zero-shot summary. CoS consistently outperforms zero-shot summarization across all models and datasets, with improvements using GPT-4o-mini as backbone on TriviaQA (0.80 vs. 0.73, a 9.6% gain) and in Llama3.2:3B on TruthfulQA (14.77 vs. 8.89, a 66% improvement). Our approach also exceeds traditional summarization methods, surpassing BRIO by 27% (0.80 vs. 0.63) and PEGASUS by 11% (0.80 vs. 0.72) on TriviaQA with GPT-4o-mini. Notably, CoS also outperforms Chain of Density, another iterative prompting approach, by 9.6% with GPT-4o-mini (0.80 vs. 0.73). This advantage becomes more pronounced with smaller models, where CoS exceeds Chain of Density by 35% with Llama3.2:3B (0.62 vs. 0.46) and by 67% with Qwen2.5:7B (0.60 vs. 0.36), suggesting that the question-guided dialectical refinement of CoS is more effective than density-based iteration, particularly when model capacity is limited. A key finding is that CoS summaries consistently present information more effectively for downstream QA purposes than the original documents themselves. For instance, on TriviaQA with a GPT-4o-mini backbone, the CoS summary achieves a score of 0.80, which is higher than the score of 0.76 achieved using the full source content. This pattern, whereby the summaries created by CoS are more useful than the raw text, is consistent across multiple models and datasets, as demonstrated in Table 1. For memorization, the difference is particularly striking with Qwen2.5:7B on TruthfulQA, where CoS achieves 46.06 compared to just 18.55 for zero-shot summaries.
@@
Figure 2. Summary length (tokens, log scale) vs. Correct-F1 score on the TriviaQA dataset, comparing performance across model backbones and summarization methods. This visualization demonstrates CoS’s ability to create information-dense summaries that optimize the efficiency-performance trade-off across model architectures.
@@
One of the questions for implementing our CoS approach is whether synthetic questions can effectively replace costly human-crafted ones. As shown in Table 3, synthetic questions are indeed viable. Across all three models, synthetic questions yielded comparable or superior results to human-crafted ones. With GPT-4o-mini, synthetic questions improved generalization from 0.73 to 0.80 (+0.07), slightly outperforming human-crafted questions, which improved from 0.74 to 0.79 (+0.05). Similarly, Qwen2.5:7B showed nearly identical improvements with synthetic questions (memorization: +0.29, generalization: +0.12) compared to human-crafted questions (memorization: +0.33, generalization: +0.10). The prompts for question generation and answering can be found in Listings 1 and 2.